​CareMore, Inc. provides​ in-home medical assistance to the elderly and earned net income of $5 million that it plans to use to repurchase shares of the​ firm's common​ stock, which is currently selling for $50 a share. CareMore has 20 million shares of stock outstanding.

a. What fraction of the​ firm's shares can the firm repurchase for $5 million?

b. If the share repurchase has no impact on the​ firm's net​ income, what will be its earnings per share after the​ repurchase?
Answer #1

net income = 5 million = 5,000,000 and price per share =50

No of shares repurchased = 5,000,000/ 50 = 100,000

(a) fraction of firm's shares repurchased = 100,000/20,000,000 = 0.5%

no of shares left after repurchase = 20,000,000 - 100,000 = 19,900,000

(b) Earnings per share after repurchase = 5,000,000/ 19,900,000 = 0.251256 = 0.25 (rounded to 2 decimals)

Thumbs up please)